A painter can finish painting a house in 3 hours. Her assistant takes 5 hours to finish the same job. How long would it take for them to complete the job if they were working together?
A. 2 hours
B. 8 hours
C. 15 hours
D. 1 7/8 hours

Answer by ewatrrr(24785)

 
Hi
Let x represent time it would take together
Question states*** Per hour being the equalizer
1/3 + 1/5 = 1/x Multiplying thru by 15x so as all denominators=1
5x + 3x = 15
8x = 15
x = 15/8 or 1 7/8 hrs, time it would take together
CHECKING our Answer***
5/15 + 3/15 = 1/(15/8) = 8/15